OpenStudy (anonymous):

If income I is $98 when 28 tickets are sold determine the income if 26 tickets are sold?

OpenStudy (anonymous):

98/28 = price of each ticket = $3.5 26*$3.5 = cost of 26 tickets $91 --------> answer. :)

Parth (parthkohli):

Income on one ticket = 3.5 dollars Income on 26 = 3.5 * 26 = 91 dollars

OpenStudy (anonymous):

Thank you both for showing your work [= so I can try to figure out how to do the next one on my own

OpenStudy (anonymous):

no prob. :)

OpenStudy (anonymous):

ParhKohli, I messaged you.